A research firm conducted a survey to determine the mean amount of money smokers spend on cigarettes during a day. A sample of 100 smokers revealed that the sample mean is N$ 5.24 and sample standard deviation is N$ 2.18 Assume that the sample was drawn from a normal population. 2.5) Use the data provided above to test a claim at a 1 % Level of Significance that the unknown population mean of spending amount on cigarettes is N$ 5.00 (8)
1
Expert's answer
2020-11-08T18:41:05-0500

The provided sample mean is "\bar{x}=5.24" and the sample standard deviation is "s=2.18," and the sample size is "n=100."

The following null and alternative hypotheses need to be tested:

"H_0:\mu=5"

"H_1:\mu\not=5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

The number of degrees of freedom are "df=100-1=99."

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a two-tailed test is "t_c=2.626405"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{5.24-5}{2.18/\sqrt{100}}\approx1.100917"

Since it is observed that "|t|=1.100917<2.626405=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu"  is different than 5, at the 0.01 significance level.


Using the P-value approach: The p-value is "p=0.2736," and since "p=0.2736>0.01," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"  is different than 5, at the 0.01 significance level.